# UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction

### v1

@@ Appendix D RQ4: Practicality
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform (Mapbox, [n. d.]). This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. Figure 10 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
@@
Each urban region captured by a satellite image, encompassing a 1km by 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.

### v2

@@ Appendix I Practicality
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform [56]. This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. Figure 17 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
@@
Each urban region captured by a satellite image, encompassing a 1km $\times$ 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.
@@
Appendix J Limitations and Social Impact
Due to the easier accessibility of the dataset, our work primarily focuses on first-tier and domestic cities, which introduces a certain bias and can be regarded as a limitation of our study. Moreover, the model’s performance is significantly influenced by image resolution; higher quality image data is required to achieve optimal performance. Additionally, the introduction of new modalities inevitably leads to a decrease in inference efficiency, which should be a consideration for practical deployment. Moreover, the inability to automatically correct low-quality text highlights an area for future improvement.
Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding urban planning and decision-making processes. However, it is crucial to note that satellite and street-view images might inadvertently contain private information. The street-view images we collected have been blurred by the Baidu Map to obscure relevant information, such as license plates, thereby significantly reducing associated risks. Nevertheless, we strongly advise users to perform a comprehensive analysis of the data prior to deploying our model in practical contexts.